How would you perform the indicated operation?
8 over the square root of 80-3 over the square root of 45

Answer by edjones(7311)

8/sqrt(80) - 3/sqrt(45)
=8/sqrt(16*5) - 3/sqrt(9*5)
=8/4sqrt(5) - 3/3sqrt(5)
=2/sqrt(5) - 1/sqrt(5)
=(2-1)/sqrt(5)
=1/sqrt(5)
=1/sqrt(5) * sqrt(5)/sqrt(5)
=sqrt(5)/5
.
Ed